A short reinforced concrete column of section 200 mm x 220mm is to be reinforced with 4 numbered steel bars and is required to carru an axial load of 850kN. The stress in the concrete must not exceed 7N/mm^2 and the stress must not exceed 150N/mm^2. Determine the diameter required for the bars and the subsequent stresses occurring in the concrete and the steel under the specified load. (Young's moduli of concrete=14kN/mm^2 and Young's moduli of steel=210kN/mm^2. With suitable sketch
1
Expert's answer
2020-02-13T06:17:31-0500

"\delta_{st}=\delta_{co}=\delta"


"(\frac{F\cdot l}{S\cdot E})_{co}=(\frac{F\cdot l}{S\cdot E})_{st}"


"\frac{\sigma_{co}\cdot l}{14\cdot10^9}=\frac{\sigma_{st}\cdot l}{210\cdot10^9}"


"14\cdot \sigma_{st}=210\cdot \sigma_{co}"


When "\sigma_{st}=150\cdot 10^6 Pa"


"14\cdot 150\cdot 10^6=210\cdot \sigma_{co} \to \sigma_{co}=10\cdot 10^6 Pa>6\cdot 10^6 Pa" (not okay!)


When "\sigma_{co}=6\cdot 10^6 Pa"


"14\cdot \sigma_{st}=210\cdot 6\cdot 10^6 \to \sigma_{st}=90\cdot 10^6 Pa<150\cdot 10^6 Pa" (okay!)


Use "\sigma_{st}=90\cdot 10^6 Pa; \sigma_{co}=6\cdot 10^6 Pa"


"F_{st}+F_{co}=850000"


"\sigma_{st}\cdot S_{st}+\sigma_{co}\cdot S_{co}=850000"


"90\cdot 10^6\cdot S_{st}+6\cdot 10^6\cdot(0.2\cdot 0.22-S_{st})=850000"


"84\cdot S_{st}=0.85-0.264 \to S_{st}=0.006976 m^2=6976m^2"


For one steel bar


"S_{0st}=6976\/4=1744 mm^2"


"\frac{\pi\cdot D^2}{4}=1744 \to D\approx 47.1 mm" Answer